# Learning from Reasoning Failures via Synthetic Data Generation

### v1

@@ 7.2 Dataset statistics
7.3 Data generation prompt
@@ 7.4 Filtering prompt
7.5 Training hyperparameters
To train our model, we used 8 Nvidia RTX A6000 GPUs using the hyperparameters from Table 6. We employed DeepSpeed ZeRO stage 3 [1] for distributed training.
@@ 7.5 Training hyperparameters
Number of GPUs $8$
Gradient Accumulation $1$
Number of epochs $1$
@@
Learning Rate $52e-$ 5
BF16 $True$
LR scheduler $cosine$
@@
Table 6: Hyperparameters to train our model.
8 Additional experimental results
8.1 Detailed OOD results for models fit to different subsets of synthetically generated data
Train Dataset $N$ $N_{syn}$ TextVQA OCR-Bench InfoVQA VQAv2 OK-VQA ScienceQA MMBench MMMU
@@ 8.1 Detailed OOD results for models fit to different subsets of synthetically generated data
Table 7: Training data augmentation experimental results. $N$ denotes the total number of examples used for training, while $N_{syn}$ denotes the number of synthetic examples in the training dataset which were generated using our approach.
@@
8.2 Generalization to other models
To evaluate the generalization ability of our synthetic dataset, we trained models using the same datasets but with different backbone LLMs. In the following experiments, we used Gemma-2B [36] and Qwen2-7B [43] as base LLMs. We adopted the LLaVA two-phase training procedure: pretraining on the LLaVA 558k dataset followed by instruction fine-tuning. Table 3 indicates that our synthetic data in its maximum augmentation setting, $N_{syn}=124,922$, outperform the baseline for both LLaVA-Gemma-2B and LLaVA-Qwen-7B almost on all benchmarks.
Model $N$ $N_{syn}$ TextVQA OCR-Bench InfoVQA VQAv2 OK-VQA ScienceQA MMBench MMMU
@@ 8.2 Generalization to other models
Table 8: Training data augmentation experimental results. $N$ denotes the total number of examples used for training, while $N_{syn}$ denotes the number of synthetic examples in the training dataset which were generated using our approach. Best result is highlighted in bold and second best result is underlined
9 Additional analyses
9.1 Detailed example comparing different LLMs used for synthetic data generation
@@
Figure 7: Comparison of GPT-4o and Qwen2-VL for generating Failure-Grounded Synthetic Datasets: GPT-4o demonstrates stronger reasoning capabilities, identifying multiple reasoning failures such as missed contextual cues and a lack of correlation between visual elements and the question. In contrast, while Qwen2-VL correctly answering the original question, identifies fewer failure modes and is less accurate in diagnosing LLaVA’s reasoning failures, sometimes focusing on less relevant aspects, such as the kite in the sky. As a result, Qwen2-VL’s generated samples are less diverse, often repeating the same question, whereas GPT-4o’s samples provide broader coverage of identified reasoning failures. Note: GPT-4o’s reasoning is 2–3 times longer than Qwen2-VL’s; only a portion of GPT-4o’s reasoning is shown here, while Qwen2-VL’s reasoning is presented in full.
10 Examples from our dataset
In this section, we present examples from our dataset and highlight its weaknesses and limitations. Figure 9 shows our synthetic data generated from the OK-VQA dataset while Figure 10 corresponds to the VizWiz dataset. In these examples, both the image, the question and the answer are all synthetic. Figure 8 provides additional fully synthetic text & image examples derived from VizWiz and OK-VQA. Figure 11, 12 and 13 illustrate examples derived from the ScienceQA, InfoVQA and VizWiz benchmarks respectively, where the images are real but the questions and answers are synthetically generated. Lastly, Figures 14 and 15 show some incorrect examples for each benchmark.

### v2

@@ A.2 Dataset statistics
A.3 Data generation prompt
@@ A.4 Filtering prompt
A.5 Training hyperparameters
To train our model, we used 8 Nvidia RTX A6000 GPUs using the hyperparameters from Table 8. We employed DeepSpeed ZeRO stage 3 (AI 2025) for distributed training.
@@ A.5 Training hyperparameters
Number of GPUs 88
Gradient Accumulation 11
Number of epochs 11
@@
Learning Rate 2e52e-5
BF16 TrueTrue
LR scheduler cosinecosine
@@
Table 8: Hyperparameters used to train models.
Appendix B Additional experimental results
B.1 Additional evaluation results for OOD datasets
We used a variety of multimodal reasoning benchmarks for evaluating models’ out-of-domain performance. Since InfoVQA and ScienceQA questions rely heavily on reading text contained in the images, we further evaluated models on TextVQA (Singh et al. 2019) and OCR-Bench (Fu et al. 2024). Finally, because our synthetic dataset was designed to enhance the model’s reasoning capabilities, we chose MMBench (Liu et al. 2024b) and MMMU (Yue et al. 2024) as additional OOD benchmarks. Evaluation results for the OOD benchmarks are provided in Table 9. While our primary focus is improving the model’s performance in-domain (i.e., correcting its reasoning failures), notably we observe that we are able to achieve this without degrading performance on these OOD tasks. This demonstrates the ability of our approach to correct reasoning failures without sacrificing the model’s general reasoning capabilities.
Base LLM NN $N_{syn}$ Aug Data T-VQA OCR-B VQAv2 MMB MMMU
Vicuna-7B 624,610 0 N/A (baseline) 47.0 31.9 71.5 52.3 36.4
Vicuna-7B 687,071 62,461 ALLaVA 47.9 34.0 72.6 50.2 36.7
CoSyn-400K 47.1 31.8 72.3 52.4 34.6
SimVQA 46.8 31.6 71.8 53.5 34.7
Ours 47.4 33.2 71.3 52.5 36.2
Vicuna-7B 749,532 124,922 ALLaVA 47.2 34.1 72.8 43.5 34.2
CoSyn-400K 46.8 32.7 72.0 51.9 36.9
Ours 47.4 34.5 72.4 52.5 37.4
Gemma-2B 624,610 0 Ours 39.9 28.3 68.7 29.2 32.3
749,532 124,922 40.9 29.9 69.0 30.7 31.7
Qwen2-7B 624,610 0 Ours 45 31.4 71.7 63 42.3
749,532 124,922 46.2 32.5 72.9 63.5 42.2
Table 9: Out-of-domain training data augmentation evaluation results. NN denotes the total number of training examples, $N_{syn}$ denotes the number of synthetic examples in the training dataset generated using our approach. T-VQA denotes TextVQA, OCR-B represents OCR-Bench, and MMB means MMBench datasets. The first section of the table compared the same base LLM (Vicuna) trained on various datasets with our dataset, while the second section compares different LLM backbones trained on our dataset.
B.2 Detailed OOD results for models fit to different subsets of synthetically generated data
Train Dataset NN $N_{syn}$ T-VQA OCR-B I-VQA VQAv2 OK-VQA SQA MMB MMMU
@@ B.2 Detailed OOD results for models fit to different subsets of synthetically generated data
Table 10: Training data augmentation experimental results. NN denotes the total number of examples used for training, while $N_{syn}$ denotes the number of synthetic examples in the training dataset which were generated using our approach. T-VQA denotes TextVQA, OCR-B represents OCR-Bench, I-VQA denotes InfoVQA, SQA represents ScienceQA and MMB means MMBench datasets.
@@
B.3 Dataset Generation Cost Comparison
For 100K synthetic samples, our pipeline required approximately 72 million GPT-4o tokens for question-answer generation and filtering, along with 418 accelerator-hours on Intel Gaudi 2 (or 766 hours on an NVIDIA A100). At current public cloud pricing, this corresponds to roughly US\$0.5K for language model API calls and US\$5.5K for Gaudi 2 compute, totaling approximately US\$6K.
In comparison, CoSyn’s code-guided renderer is dominated by text-only LLM calls, as reported in their paper, generating 100K samples costs approximately US\$2K. The relatively low computational cost of generating chart- and OCR-style images makes CoSyn a natural complement to our method: CoSyn provides structured, code-generated visuals, while our pipeline contributes rich, multimodal failure-supervised content. ALLaVA uses a slightly simpler caption + question + answer format than ours, omitting our failure-mode identification step. However it provides GPT-4V with high-resolution real images, whereas we downsample to $125\times 125$ pixels to reduce input token count. Additionally, ALLaVA exclusively uses real images, while we optionally generate synthetic images, adding compute cost in exchange for controllability and diversity. SimVQA’s Hypersim/TDW pipeline incurs significantly higher generation cost due to photorealistic 3D rendering on GPU clusters. This includes simulating realistic and dynamic environments, making it highly compute-intensive and resource-heavy. Finally Img-Diff, which generates pairs of images that are visually alike but differ in specific objects. They first generate similar images where some objects are replaced using an LMM and a diffusion model (SDXL (Podell et al. 2023) and Prompt-to-prompt (Hertz et al. 2022)). Then, they identify the locations of object differences. They extract bounding box regions that contain object differences using CLIP (Radford et al. 2021) to extract the degree of similiraty, then they use FastSAM (Zhao et al. 2023) to perform image segmentation on each image. Based on the segmentation, they crop the images and use BLIP (Li et al. 2022) to determine whether an image contains valid objects. Lastly, they use a MLLM to generate descriptive text for the areas with object differences and creates question-answer pairs. In order to run their pipeline and obtain 118k high-quality image pairs, it took four NVIDIA A100 GPUs for 4.5 days.
Overall, our dataset offers a balanced compromise between cost, quality, and diversity, and can be effectively combined with lower-cost image generation pipelines such as CoSyn to enable scalable and semantically rich supervision
B.4 Generalization to other models
To evaluate the generalization ability of our synthetic dataset, we trained models using the same datasets but with different backbone LLMs. In the following experiments, we used Gemma-2B (Team et al. 2024) and Qwen2-7B (Yang et al. 2024) as base LLMs. We adopted the LLaVA two-phase training procedure: pretraining on the LLaVA 558k dataset followed by instruction fine-tuning. Table 2 indicates that our synthetic data in its maximum augmentation setting, $N_{syn}=124,922$, outperform the baseline for both LLaVA-Gemma-2B and LLaVA-Qwen-7B almost on all benchmarks.
Model NN $N_{syn}$ T-VQA OCR-B I-VQA VQAv2 OK-VQA SQA MMB MMMU
@@ B.4 Generalization to other models
Table 11: Training data augmentation experimental results. NN denotes the total number of examples used for training, while $N_{syn}$ denotes the number of synthetic examples in the training dataset which were generated using our approach. T-VQA denotes TextVQA, OCR-B represents OCR-Bench, I-VQA denotes InfoVQA, SQA represents ScienceQA and MMB means MMBench datasets. Best result is highlighted in bold and second best result is underlined
Appendix C Additional analyses
C.1 Detailed example comparing different LLMs used for synthetic data generation
@@
Figure 6: Comparison of GPT-4o and Qwen2-VL for generating Failure-Grounded Synthetic Datasets: GPT-4o demonstrates stronger reasoning capabilities, identifying multiple reasoning failures such as missed contextual cues and a lack of correlation between visual elements and the question. Qwen2-VL also correctly answering the original question, but identifies fewer failure modes. Note: GPT-4o’s reasoning is 2–3 times longer than Qwen2-VL’s; only a portion of GPT-4o’s reasoning is shown here, while Qwen2-VL’s reasoning is presented in full.
Appendix D Examples from our dataset
In this section, we present examples from our dataset and highlight its weaknesses and limitations. Figure 7 shows illustrative examples from the OK-VQA dataset. For both samples, we show the original image, question, and answer, followed by the failure mode identified by GPT-4o. We also include newly generated questions, images, and ground truth answers designed to reflect the same failure mechanism, as well as the answer predicted by our finetuned model (on +124k samples).
Figure 7: OK-VQA examples with original image-question-answer pairs, the reasoning-based failure identified by GPT-4o, and newly generated counterparts.
Figure 9 shows our synthetic data generated from the OK-VQA dataset while Figure 10 corresponds to the VizWiz dataset. In these examples, both the image, the question and the answer are all synthetic. Figure 8 provides additional fully synthetic text & image examples derived from VizWiz and OK-VQA. Figure 11, 12 and 13 illustrate examples derived from the ScienceQA, InfoVQA and VizWiz benchmarks respectively, where the images are real but the questions and answers are synthetically generated. Lastly, Figures 14 and 15 show some incorrect examples for each benchmark.
@@ Appendix D Examples from our dataset
D.1 Impact of filtering on data quality
D.2 Training data substitution vs. augmentation and impact of synthetically generated images
Train Data NN $N_{syn}$ T-VQA OCR-B I-VQA VQAv2 OK-VQA SQA MMB MMMU
Baseline 624,610 0 47.0 31.9 26.7 71.5 57.0 70.7 52.3 36.4
624,610 156153 46.9 31.1 27.0 72.7 57.0 70.6 51.2 37.9
687,071 62,461 47.7 31.8 25.8 71.6 59.4 71.2 52.3 36.4
687,071 62,461 47.2 32.2 27.2 71.8 56.9 71.2 52.3 33.8
Table 12: Ablation experiments comparing baseline LLaVA to LLaVA models trained with synthetic data generated from VizWiz failures. T-VQA denotes TextVQA, OCR-B represents OCR-Bench, I-VQA denotes InfoVQA, SQA represents ScienceQA and MMB means MMBench datasets. We investigate substitution and augmentation strategies for synthetic data, as well as the use of synthetic vs. real images.
Our previous experiments augmented an existing 624k sample training dataset (LLaVA-Instruct) with our synthetic data. In domains where data is scarce, training datasets of this size may not be available. To investigate the utility of our synthetic data in such low-resource settings, we conducted experiments in which we randomly substituted different quantities of examples from the original dataset with our synthetically generated data [^2] We used synthetic data derived from Vizwiz failures in this setting.. The results of this experiment are provided in rows 2-3 of Table 12. Even when up to 25% of the original dataset is substituted with our synthetic data, we achieve performance that is either as good or better than the baseline LLaVA model across a broad range of downstream tasks. This is despite the fact that the original LLaVA training dataset utilizes real images, whereas our synthetic data used in this experiment contained only synthetically generated images. The fact that our synthetic data achieves similar or better performance than an existing real data source is significant, as prior studies have shown that training on synthetically generated image data is often much less efficient than training on an equivalent amount of real image data (He et al. 2022). Table 12 also shows the impact of using real vs. synthetic images in our pipeline. Specifically, we compare the effectiveness of our synthetic data derived from Vizwiz reasoning failures when paired with real images (from Vizwiz) or synthetically generated images. In the training data augmentation setting, we observe that synthetic images generally achieve similar results as utilizing real images. Synthetic images even surpass the performance of real images in TextVQA, OK-VQA, and MMBench. This demonstrates the high quality of our synthetic images and their potential to serve as replacements for real images in low-resource settings where data is scarce.